Can someone help me solve this problem?
I need to evaluate the following problem and need to use the order of operations:
4 - (-3) + 1/2-2 divided by (-2)
Thanks,
Ashley

Answer by tutorcecilia(2144)


You will need to apply the order of operations which is the order in which a problem is solved. The abbreviation P.E.M.D.A.S. will give you the correct order to follow:
P = Parenthesis first
E = Exponents or radicals next
M or D = Multiplication or Division (whichever comes first)
A or S = Addition or Subtraction (whichever comes first)
.
Starting from left to right, you should see that this problem only involves three operations: subtraction, addition, and division. To make it easier, convert the fraction to a decimal.:
4- (-3) + 0.50 - 2/-2
.
According to PEMDAS, perform addition or subtraction in the order in which it occurs. The subtraction was changed to addition using the rules for adding or subtracting integers*:
4+3+0.50-2/-2
7+.50-2/-2
7.50-2/-2
5.50/-2
.
Next, perform multiplication or division which ever occurs first. Since there is no multiplication, divide:
5.50/-2
-2.75
.
*Rules for adding or subtracting integers:
When adding integers with the same sign (both negative or both positive), add the absolute values and take the common sign.
When adding integers with different signs (one is positive and the other is negative), subtract the absolute values and take the sign of the larger integer.
When subtracting, change the operation to addition and change the sign of the second integer. Than follow the rules of adding integers.
Example: 4-(-3)= 4+ (+3) = +7